A ball is thrown downward with a velocity of 12.0 m/s from a height of 92.2 meters. What is the ball’s velocity when it hits the ground? How long was the ball in the air?


1
Expert's answer
2020-09-21T08:13:33-0400

The ball’s velocity when it hits the ground

"v=\sqrt{v_0^2+2gh}\\
=\sqrt{12.0^2+2\times 9.81\times 92.2}=44.2\:\rm m/s"

The time of motion

"t=\frac{h}{v_{\rm ave}}=\frac{2h}{v_0+v}""=\frac{2\times 92.2}{12.0+44.2}=3.28\:\rm s"